Documenta Mathematica

Band 15, 2010

Jianqiang Zhao Standard Relations of Multiple Polylogarithm

Values at Roots of Unity  1–34

Indranil Biswas and Norbert Hoffmann The Line Bundles on Moduli Stacks

of Principal Bundles on a Curve  35–72

José I. Burgos Gil and Răzvan Liţcanu Singular Bott-Chern Classes and the Arithmetic Grothendieck

Riemann Roch Theorem for Closed Immersions  73–176

Vanessa Miemietz and Will Turner

Hicas of Length ≤ 4  177–205

Oliver Matte On Higher Order Estimates

in Quantum Electrodynamics  207–234

iv